### v2

# Doping of a Borexino-like Liquid Scintillator with Tellurium-Diols

Abstract
One of the most promising approaches for future neutrinoless double beta decay ($0\nu\beta\beta$) searches is to incorporate a candidate isotope into the liquid scintillator of a next-generation neutrino detector. In this study, a sample of the high-performance 1,2,4-Trimethylbenzene-based liquid scintillator from the Borexino detector was loaded with different concentrations of Te-diols. Therefore, a novel and completely water-free synthesis in a non-acidic organic environment at room temperature was used. Key parameters of the loaded samples were analyzed and compared with those of the pure Borexino liquid. Both the emission spectrum and transmission remained nearly unchanged, even at high doping levels. The reduction in light yield was moderate, with approximately 8,400 photons emitted for a 1 MeV energy deposition by an electron at 1% tellurium loading. The time profile of the light emission induced by alpha particles was also investigated, revealing that the scintillation response becomes significantly faster with increasing tellurium concentration.
1 Introduction
The quest to detect neutrinoless double beta decay ($0\nu\beta\beta$) is a top priority in contemporary particle physics. Discovering this process, which violates lepton-number conservation, would not only reveal the absolute neutrino mass scale but also provide valuable insights into Grand Unified Theories and leptogenesis in the early universe  [1]. To detect neutrinoless double beta decay, several experimental technologies have been developed over the last decades. The core idea is to incorporate into the detector’s active mass a ($0\nu\beta\beta$) candidate isotope: high-purity germanium detectors (LEGEND: 76Ge), loaded scintillators (KamLAND-Zen: 136Xe, SNO+: 130Te), and xenon-based time projection chambers in liquid or high-pressure gas form (EXO and NEXT: 136Xe) are among the current experiments. In particular, liquid scintillator (LS) detectors are highly sensitive instruments in the search for this ultra-rare process  [2, 3]. The detection method offers several advantages, including extremely low background radiation, flexibility in detector shape and size, and the ability to incorporate a large quantity of the $\beta\beta$-decaying isotope directly into the LS. The SNO+ collaboration has conducted successful R&D toward realizing a liquid scintillator loaded with compounds containing natural tellurium. The cocktail uses the widely employed organic solvent LAB (linear alkylbenzene) and the fluor PPO (2,5-diphenyloxazole) [4, 5, 6, 7], while the dopants are oligomers of Te-diols. These were prepared from an aqueous telluric acid solution with the addition of diols (typically 1,2-butanediol) and stabilized with DDA (N,N-dimethyldodecylamine) [7].
2 Te-diol Synthesis and Doping 1,2,4-Trimethylbenzene-based Liquid Scintillator
For the work described here, the authors were provided with a generous amount of scintillator from the Borexino detector [8] by the Borexino Collaboration. The sample was carefeully stored under protective nitrogen atmosphere in glass bottles and subsequently loaded with Te-diols in the laboratory without any modification. The scintillator is based on the widely-used solvent 1,2,4-Trimethylbenzene (CAS: 95-63-6) known for its excellent energy transfer efficiency and 1.5 g/l 2,5-Diphenyloxazole (CAS: 92-71-7) as fluor [8].
To realize the loading of the 1,2,4-Trimethylbenzene-based LS the techniques described in [7] were reviewed and modified towards a completely water-free synthesis in a non-acidic organic environment. The process enables a Te-diol synthesis at room temperature thus not requiring heating the reactants. An only partially comparable approach was chosen by Ding et al for [9, 10]. For the synthesis of the Te-diols, reported in the present publication, the chemicals as listed in Table 1 were purchased in the highest purity directly available from Sigma-Aldrich and not further purified.
Substance Linear Formula CAS Number Purity Supplier
Orthotelluric acid (TeAs) H6TeO6 7803-68-1 purum, $\geq$99.0% Sigma-Aldrich
1,2-Butanediol (BD) CH3CH2CH(OH)CH2OH 584-03-2 purum, $\geq$98.0% Sigma-Aldrich
1-(Dimethylamino)dodecane (DDA) CH3(CH2)11N(CH3)2 112-18-5 $\geq$97.0% Sigma-Aldrich
Benzene C6H6 71-43-2 for HPLC, $\geq$99.9% Sigma-Aldrich
Table 1: Shown are the chemicals used for the synthesis of the Te-diol compounds. The table lists the linar furmulae as well as CAS numbers. Note that all Chemicals were purchased as pure as directly available from Sigma-Aldrich and therefore not futher purified before the synthesis.
For the synthesis, the orthotelluric acid was first crushed in a mortar to obtain a powder with a fine grain size. The telluric acid was then given into a round flask and neutralized with 1-(dimethylamino)dodecane at room temperature while stirring constantly. Benzene is then added to the mixture. To form the Te-diol complexes in the liquid 1,2-Butanediol is now slowly added using a burette. During this process, water is produced in the reaction, which is removed by constant bubbling the liquid with dried nitrogen. Typically within minutes the milky and cloudy liquid clears up and gets fully colorless and transparent while stirring continuously. For the synthesis the molar ratios of the compounds (similarly to [7]) are 0.5:1:2 (DDA:TeA:BD). A tiny excess of a few milligrams of TeAs ensures that no BD remains, which would have a negative effect on the transparency of the final scintillator, as it is almost insoluble in PC. To remove potential TeAs crystals from the liquid filtration by means of a Büchner funnel equipped with ash-free filter paper (retention: 2 µm) is performed. However, it should be mentioned that only negligible amounts of telluric acid remain in the filter paper. In the final step, the benzene is evaporated at room temperature with constant stirring, leaving behind a colorless, transparent, and highly viscous liquid.
It should be noted here that benzene was deliberately chosen to enable evaporation at room temperature and normal pressure of the organic carrier medium. This ensures the removal of the by-product of the reaction (water) by means of an azeotropic distillation in a simple manner. It is clear that benzene could be replaced by toluene or cyclohexane if necessary, and that azeotropic distillation could be carried out using a rotary evaporator and low pressure at similarly low temperatures. Despite its toxicity, benzene has proven to be ideal and particularly easy to use for the small quantities of Te-diols required here, while toluene should be preferred for later, larger applications.
To load the 1,2,4-Trimethylbenzene-based LS it is added to the Te-diols in a quantity that ensures a loading of 2% (by mass) of Tellurium in final liquid. Note that the mass-loading fraction with the Te-diol product is higher by a factor of about 2.67. Typically the dopant dissolves in the LS completely within seconds. The lower concentrations also described in the later chapters of this publication are produced by dilution of this concentrated 2% mixture with additional Borexino scintillator. All scintillators are extensively flushed with dried nitrogen gas and stored in light-tight glass bottles under protective atmosphere.
3 Characterization of Fundamental Scintillator Properties
In recent decades, the requirements for scintillators in the field of neutrino physics have become increasingly demanding. Of particular importance are high transmission for the light emitted by the scintillator, particularly high light yields with spectra matching well the spectral response function of conventional photomultipliers, and a well-characterized and sufficiently fast time profile of light emission. The pseudocumene-based scintillator used in the Borexino experiment has set new standards in this regard, in addition to its outstanding radio purity [8, 11]. In this section, the Borexino sample is compared with regard to the properties of spectral light emission, spectral absorbance, light yield, and time profile for alpha particle interaction with mixtures of varying Te-loading in order to evaluate the influence of increasing amounts of Te-diols on these key properties of the scintillator.
3.1 Light Emission Spectra
To evaluate the effective emission spectrum of the scintillator mixtures (as seen by photosensors in a detector), the samples with different amounts of Te-diols were examined and compared with that of the unloaded mixture of PC and PPO using an Edinburgh FS5 spectro-fluorometer. To take into account the self-absorption of the samples, they were measured pure and undiluted in a $10\times 10\times 40$ mm3 fused silica cuvette. The conventional geometry sample holder (Edinburgh SC-05) ensures a long light path through the sample. All scintillators were excited with light of (260 $\pm$ 1) nm close to the PC’s maximum absorption at 267 nm, where 0.5 nm was selected as emission and 1 nm excitation slit-widths. All measured light emission spectra are identically dominated by the typical emission profile of PPO with a clear maximum at $\sim$365 nm and a pronounced tail towards higher wavelength. No influence on the shape of the effective emission spectrum of the light could be found. For clarity, the measurements for the unloaded Borexino scintillator and a sample with 1% Te-loading are shown in Figure 1. However, it should be noted here that although the spectral shape is not changed, the loading does have an influence on the amount of emitted light. The studies for this influence on the global light yield are presented in subsection 3.3.
Figure 1: Effective emission spectra for the unloaded Borexino scintillator and a sample loaded with 1% Te. Within the measurement accuracy, no differences in spectral shape were detected. Both spectra are completely dominated by the fluor PPO.
3.2 Spectral Transmission of Light
In order to evaluate the transparency and thus a potential reduction due to increased Te-doping, the absorbance of the undoped scintillator was compared with the loaded samples. Therefore, a Perkin Elmer Lambda 850+ UV/Vis-spectrometer with a long rectangular fused-silica cuvette (10 cm optical path length) was used to gain the results shown in Figure 2.
Within the scope of measurement uncertainties, no relevant deterioration in the absorbance of the samples could be observed. Loading with 0.125 % - 2 % of the Te-complex appears to have no relevant influence on the transmission of the Borexino mixture. For reasons of clarity, only the pure scintillator and the sample loaded with 1 % are shown in Figure 2.
Figure 2: Comparative plot of the spectral absorbance of undoped and 1 % Te-doped LS. The bands shown, represent the statistical uncertainty from ten individual measurements in each case. Due to the relatively short cuvette length, it should be noted that the systematic uncertainties are significant but are not shown for the sake of clarity.
3.3 Light Yields
The amount of scintillation photons produced for a specific energy deposition by a specific particle species, the LS light yield (LY), is a crucial parameter greatly affecting the energy resolution and vertex reconstruction capabilities of a scintillation medium. In a two component scintillator (solvent + fluor) the LY drastically increases with increasing fluor loadings from approximately 1 g/l - 10 g/l. Here, internal losses in the solvent (mainly due to impurities) compete with the energy transfer from the solvent to the fluor. At the critical concentration (typ. <1<1 g/l) for the solvent-fluor combination, about half the maximum light yield is reached [12, 13]. For scintillation cocktails with effective energy transfers between solvent and fluor the critical concentration is lower, leading to higher light yields with lower amounts of fluor.
Absolute measurements of the amount of scintillation photons emitted for a given energy deposition in a scintillator liquid are very difficult and prone to mismeasurements. Therefore, most light yields are determined relative to a reference scintillator with a known LY [14]. For the LY studies presented here, this method was also applied with the undoped Borexino LS as a reference.
To evaluate the light yield of the liquid scintillators, the setup shown in Figure 3 and described in detail in [15] was used. A 10.9 ml sample of LS is filled under protective N2 atmosphere into a highly reflective PTFE cell with 2 mm thin UV-transparent glass windows such that the entire volume is free of gas bubbles. Before and after the filling the sample as saturated with protective inert gas by extensively bubbling it with N2. To protect the sample, the cell is sealed under moderate over-pressure. For the readout of the scintillation light, two conventional 1.13-inch PMTs (ETEL9128B) are coupled to the cell windows. During the measurements the sample is irradiated by mono-energetic 662 keV gammas from a 137Cs source (activity: $\sim$370 kBq). The position of the source is thereby fixed with respect to the LS cell. The setup is triggered on a coincidence with a 1.5 inch $\times$ 1.5 inch LaBr3(Ce) detector provided by OST Photonics with 3.05% resolution at 662 keV. In the offline analysis only events with a coincident signal in the 185 keV backscattering pseudo-peak of the crystal detector and thus a 477 keV energy deposition in the liquid scintillator are selected [15, 16].
The measured values for different loadings are shown in Table 2 and graphically depicted in Figure 4. By studying the reproducibility of the light yield measurements for the samples the combined systematic uncertainties were estimated. Each sample was measured 10 times on different days each time freshly filled into the previously cleaned empty cell. The variance of these measurements is interpreted as the influence of systematic effects like HV stability, minor temperature changes in the laboratory, efficiency of the oxygen removal from the LS and other sample preparation and cleaning related differences. By this study the relative systematic uncertainty on the LY of a given sample can be estimated as well as statistical errors.
Figure 3: Schematic drawing of the LY setup. A small $1\times 1$ inch highly reflective PTFE liquid scintillator cell with thin UV-transparent glass windows is filled with the LS sample. Two PMTs are coupled to the cell. The LS is irradiated by monoenergetic gamma-quanta from a 137Cs source. To fix the scattering angle and by that the energy deposition of the gammas in the LS, the setup is triggered in coincidence with a LaBr3(Ce) detector. For further details on the setup and the measurement technique see also [15].
Te Loading 0 % 0.125 % 0.25 % 0.5 % 1.0 % 2.0 %
Rel. LY [%] 100 $91.5\pm 3.42$ $82.3\pm 3.63$ $71.1\pm 3.50$ $61.9\pm 3.95$ $41.8\pm 4.01$
Abs. LY [Photons/MeVee] 13600 $2444\pm 466 11$$1193\pm 494$ $9670\pm 476$ $8418\pm 538$ $5685\pm 546$
Table 2: The table shows the relative light yield values together with the Te concentration in the scintillator. The last row shows absolute values for the LY calculated using data provided by the Borexino Collaboration for energy depositions by electrons in their scintillator [17].
Figure 4: The plot shows the relative light yield of the sample vs. its Te-concentration. It should be mentioned, that light yields above 104 Photons for electron-like energy depositions are still realized with loadings well exceeding 0.25 %. For loadings exceeding 1.0 % the chemical quenching by then Te-complex molecules severely reduces the light emission in this simple LS baesed on a single solvent (PC) and a single fluor (PPO).
3.4 Light Emission Time Profile
The pulses recorded by photomultiplier tubes (PMTs) of a scintillation detector are shaped not only by the detector hardware but also, more fundamentally, by the intrinsic fluorescence time profile of the scintillator itself. This time profile is determined by both the composition of the scintillator and the differential energy loss $dE/dx$ of particles passing through it. For theoretical background and additional empirical data, see references [18, 19, 20]. Therefore, it is essential to evaluate the effect of increased Te-doping on the light emission time profile of the respective liquid scintillation cocktail.
3.4.1 Light emission time profile setup
To determine the light emission time profile, the setup for the characterization of the JUNO LS at the University of Milan was re-used [21]. It is composed of an UV-transparent glass cell of $\sim$ 3x3x3 cm3 in which the liquid scintillator is filled. Radioactive sources can be deployed by screwing them into the vessel, as shown in left panel of Figure 5. For all the measurements the cuvette was flushed with Nitrogen for more than 1h after filling to avoid potential oxygen-caused quenching of the time profiles. Furthermore, a slight overpressure of the N2 ensures a protective inert atmosphere in the vessel during the measurements.
Two PMTs are used to read out the scintillation light of the particle interactions. The first one is strongly coupled with the cuvette (called High Level PMT), while the second one (called Low Level PMT) is weakly coupled, by using a neutral density (ND) filter, allowing single photon counting. A detailed image of the PMT arrangement is shown in the right panel of Figure 5. The combination of a strongly and a weakly coupled PMT allows the usage of the well-established time correlated single photon counting (TCSPC) technique [22, 23, 24]. To ensure that the mandatory single photon condition is satisfied during the measurements, the charge spectrum of the weakly coupled PMT is acquired [21].
Figure 5: Left: Cuvette filled with liquid scintillator before being flushed with nitrogen. At the top center of the cuvette it is possible to see the 244Cm source screwed to the cup. Right: Time profile experimental. In is possible to see: the cuvette on the bottom left, with inserted an optical fiber for the PMT characterization, the HL-PMt on the top left very close to the cell and on the right a neutral optical filter, behind which the $L$-PMT is mounted.
The time response of the setup (IRF: Instrumental Response Function) is characterized in situ by the usage of a diode laser (EPL-405) with a typical pulse width of $\sim$ 70 ps. For the PMTs used in the setup the sharp light pulse can be treated as a delta function and thus enabling a direct characterization of the instrumental response of the LS detector (for further details see [21]). The setup reaches an intrinsic time resolution for the TCSPC of $\sim 0.45$ ns [21].
To further enhance the purity of the time profile samples, the experimental setup is placed inside a muon veto system, composed of two 500 mm × 500 mm × 20 mm EJ-200 plastic scintillator detectors.
3.4.2 Time Profiles for $\alpha$-Interactions with the Liquid Scintillator
Figure 6: Fluorescence time profile of the different mixtures normalised at maximum. On the left, the whole profile is shown, while on the right, a zoom in the interested region has been plotted. Due to the quenching effect on the light yield, the statistics for high concentrated Tellurium mixture is lower but enough to describe the fluorescence profile in the first hundreds of nanoseconds.
For the time profile study presented in this publication a 244Cm alpha source was deployed into the LS vessel. After extensive flushing with nitrogen gas the data taking took place subsequently. The time profiles for different tellurium loadings of the LS are depicted in Figure 6 in logarithmic scale for both, the full DAQ window of 1600 ns (left plot) and a zoom in (right plot) on the first $\sim$ 100 ns of the light emission decay process.
To better quantify changes in the time profile, a fit model based on four exponential decay functions convoluted with the IRF is applied to every dataset. The model can be denoted as equation 3.1. A more detailed description of the setup and the model are given in [21].
$F_{fit}(t)=IRF(t-t_{0})\otimes N\sum_{i=1}^{4}\frac{A_{i}}{\tau_{i}-\tau_{r}}\left(e^{-\frac{t-t_{0}}{\tau_{i}}}-e^{-\frac{t-t_{0}}{\tau_{r}}}\right)$ (3.1)
The results of all the fits are shown in table 3, in which it is possible to see the reduction of the decay time components for the all the components, while the fractions A are almost constant.
Te Loading 0 % 0.125 % 0.25 % 0.5 % 1.0 % 2.0 %
$A_{1}$ (%) $68.15\pm 0.15$ $9.17\pm 0.16 66$$6.75\pm 0.31$ $67.47\pm 0.19$ $8.10\pm 0.43 66$$6.86\pm 0.43$
$A_{2}$ (%) $19\pm 0.18 14.14$$14\pm 0.14$ $14.68\pm 0.27$ $14.44\pm 0.21$ $14.68\pm 0.30$ $17.30\pm 0.30$
$A_{3}$ (%) $9.60\pm 0.16$ $9.76\pm 0.11$ $10.12\pm 0.23$ $9.76\pm 0.17$ $9.27\pm 0.26$ $8.79\pm 0.26$
$A_{4}$ (%) $8.20\pm 0.29$ $7.20\pm 0.24$ $8.80\pm 0.48$ $8.60\pm 0.33$ $8.00\pm 0.59$ $7.20\pm 0.59$
$\tau_{1}$ (ns) $3.212\pm 0.001$ $2.789\pm 0.004$ $2.549\pm 0.008$ $2.341\pm 0.001$ $2.010\pm 0.001$ $1.501\pm 0.001$
$\tau_{2}$ (ns) $15.05\pm 0.29$ $2.59\pm 0.20 11$$38\pm 0.35 11.$$1.48\pm 0.25$ $9.62\pm 0.25$ $6.59\pm 0.24$
$\tau_{3}$ (ns) $74.10\pm 2.70$ $67.57\pm 1.12$ $65.36\pm 3.07$ $61.30\pm 2.40$ $55.90\pm 3.20$ $46.92\pm 3.20$
$\tau_{4}$ (ns) $478.9\pm 30.3$ $470.0\pm 30.0$ $82.7\pm 40.2 44$$48.9\pm 30.4$ $385.0\pm 51.6$ $399.8\pm 51.6$
Table 3: Parameters of the $\alpha$ time profile models for different Te-loadings. Provided values include statistical and systematic uncertainties. Increased loading leads to significantly faster scintillation light decay.
4 Conclusions
Metal-loaded scintillators will play an important role in the search for neutrino-less double beta decay in the future. Experiments such as SNO+ and KamLAND-Zen have already achieved initial success in this area. In this work, a sample of the scintillator from the Borexino experiment was successfully loaded with Te-diols. To realize this, the techniques of the SNO+ experiment were reviewed and modified towards a completely water-free synthesis in a non-acidic organic environment. The process enables a Te-diol synthesis at room temperature thus not requiring heating the reactants. Key parameters of the loaded sample were examined and compared with the initial sample. The emission spectrum and transmission were virtually unaffected by even high doping factors. The decline in light yield is also moderate, meaning that even with 1% tellurium, approximately 8400 photons for 1 MeV energy deposition by an electron are emitted. The time profile of light emission by alpha particles was also studied. It was shown that the speed of scintillation increases significantly when the tellurium loading is increased. This effect was expected, as the addition of large atoms or molecules to scintillators usually leads to chemical quenching effects that accelerate light emission while reducing the absolute light yield. Overall, it was demonstrated that the loading technique used here is suitable even for pseudocumene-based scintillators, which are known to be very sensitive.
Acknowledgments
This work benefited substantially from the support with seed funding from the Cluster of Excellence ORIGINS (DFG, German Research Foundation under Germany’s Excellence Strategy – EXC-2094 – 390783311).
This work benefited from the support by the Detector Laboratory of the Cluster of Excellence PRISMA+. We are very grateful to the staff of the Detector Division around Dr. Quirin Weitzel. We would like to express our warmest thanks to Dr. Mihail Mondeshki (Chemistry Department, JGU Mainz) and his team for their support with chemical expertise, their unbiased interdisciplinary collaboration, and numerous discussions.
We are grateful to the University of Milan and to INFN of Milan which host the fluorescence time profile setup in the Liquid Scintillator laboratory. We would also thank Dr. Barbara Cacciania and Fatima Horuia for the help for this work.
For countless detailed and inspiring discussions we would like to thank especially Prof. Dr. Franz von Feilitzsch (TUM), Dr. Gioacchino Ranucci (INFN Milano), Prof. Dr. Caren Hagner (University of Hamburg and DESY), Prof. Dr. Gabriel Orebi-Gann (UC Berkeley and LBNL) and Dr. Minfang Yeh (BNL). Moreover, we would like to thank Prof. Dr. Peter Fierlinger (TUM) and PD Dr. habil. Joachim Diener (TUM) for the support of the work carried out in Munich.

### v3

@@
 Therefore, a novel and completely water-free synthesis in a non-acidic organic environment at room temperature was used.
-Key parameters of the loaded samples were analyzed and compared with those of the pure Borexino liquid.
-Both the emission spectrum and transmission remained nearly unchanged, even at high doping levels.
-The reduction in light yield was moderate, with approximately 8,400 photons emitted for a 1 MeV energy deposition by an electron at 1% tellurium loading.
+Key parameters of the loaded samples were analyzed and compared with those of the pure Borexino liquid scintillator.
+Both the emission spectrum and optical transmission remained nearly unchanged, even at high doping levels.
+The reduction in light yield was moderate: approximately 8,400 photons were emitted for a 1 MeV energy deposition by an electron at 1% tellurium-loading.
 The time profile of the light emission induced by alpha particles was also investigated, revealing that the scintillation response becomes significantly faster with increasing tellurium concentration.
@@
 These were prepared from an aqueous telluric acid solution with the addition of diols (typically 1,2-Butanediol) and stabilized with DDA (N,N-Dimethyldodecylamine) [7].
-2 Te-diol Synthesis and Doping 1,2,4-Trimethylbenzene-based Liquid Scintillator
+2 Te-diol Synthesis and Doping of 1,2,4-Trimethylbenzene-based Liquid Scintillator
 For the work described here, the authors were provided with a generous amount of scintillator from the Borexino detector [8] by the Borexino Collaboration.
-The sample was carefeully stored under protective nitrogen atmosphere in glass bottles and subsequently loaded with Te-diols in the laboratory without any modification.
+The sample was carefully stored under protective nitrogen atmosphere in glass bottles and subsequently loaded with Te-diols in the laboratory without any modification.
 The scintillator is based on the widely-used solvent 1,2,4-Trimethylbenzene (CAS: 95-63-6) known for its excellent energy transfer efficiency and 1.5 g/l 2,5-Diphenyloxazole (CAS: 92-71-7) as fluor [8].
-To realize the loading of the 1,2,4-Trimethylbenzene-based LS the techniques described in [7] were reviewed and modified towards a completely water-free synthesis in a non-acidic organic environment.
+To realize the loading of the 1,2,4-Trimethylbenzene-based LS, the techniques described in [7] were reviewed and modified towards a completely water-free synthesis in a non-acidic organic environment.
 The process enables a Te-diol synthesis at room temperature thus not requiring heating the reactants.
-An only partially comparable approach was chosen by Ding et al for [9, 10].
+A partially comparable approach was reported by Ding et al.
+for [9, 10].
 For the synthesis of the Te-diols, reported in the present publication, the chemicals as listed in Table 1 were purchased in the highest purity directly available from Sigma-Aldrich and not further purified.
 Substance Linear Formula CAS Number Purity Supplier
-Orthotelluric acid (TeAs) H6TeO6 7803-68-1 purum, $\geq$99.0% Sigma-Aldrich
+Orthotelluric acid (TeA) H6TeO6 7803-68-1 purum, $\geq$99.0% Sigma-Aldrich
 1,2-Butanediol (BD) CH3CH2CH(OH)CH2OH 584-03-2 purum, $\geq$98.0% Sigma-Aldrich
-1-(Dimethylamino)dodecane (DDA) CH3(CH2)11N(CH3)2 112-18-5 $\geq$97.0% Sigma-Aldrich
+N,N-Dimethyldodecylamine (DDA) CH3(CH2)11N(CH3)2 112-18-5 $\geq$97.0% Sigma-Aldrich
 Benzene C6H6 71-43-2 for HPLC, $\geq$99.9% Sigma-Aldrich
 Table 1: Shown are the chemicals used for the synthesis of the Te-diol compounds.
-The table lists the linar furmulae as well as CAS numbers.
-Note that all Chemicals were purchased as pure as directly available from Sigma-Aldrich and therefore not futher purified before the synthesis.
-For the synthesis, the orthotelluric acid was first crushed in a mortar to obtain a powder with a fine grain size.
-The telluric acid was then given into a round flask and neutralized with 1-(dimethylamino)dodecane at room temperature while stirring constantly.
-Benzene is then added to the mixture.
-To form the Te-diol complexes in the liquid 1,2-Butanediol is now slowly added using a burette.
-During this process, water is produced in the reaction, which is removed by constant bubbling the liquid with dried nitrogen.
-Typically within minutes the milky and cloudy liquid clears up and gets fully colorless and transparent while stirring continuously.
-For the synthesis the molar ratios of the compounds (similarly to [7]) are 0.5:1:2 (DDA:TeA:BD).
-A tiny excess of a few milligrams of TeAs ensures that no BD remains, which would have a negative effect on the transparency of the final scintillator, as it is almost insoluble in PC.
-To remove potential TeAs crystals from the liquid filtration by means of a Büchner funnel equipped with ash-free filter paper (retention: 2 µm) is performed.
+The table lists the linear formulae as well as CAS numbers.
+Note that all chemicals were purchased as pure as directly available from Sigma-Aldrich and therefore not further purified before the synthesis.
+For the synthesis, the orthotelluric acid was first ground in a mortar to obtain a fine powder.
+The telluric acid was then added to a round flask and neutralized with N,N-Dimethyldodecylamine at room temperature while stirring constantly.
+Benzene was then added to the mixture.
+To form the Te-diol complexes in the liquid 1,2-Butanediol was slowly added using a burette.
+During this process, water is produced in the reaction, which is removed by continuously bubbling dried nitrogen through the liquid.
+Typically within hours the milky and cloudy liquid clears up and becomes fully colorless and transparent while stirring continuously.
+For the synthesis, the molar ratios of the compounds (similarly to [7]) are 0.5:1:2 (DDA:TeA:BD).
+A tiny excess of a few milligrams of TeA ensures that no BD remains, which would have a negative effect on the transparency of the final scintillator, as it is almost insoluble in PC.
+To remove potential TeA crystals from the liquid filtration by means of a Büchner funnel equipped with ash-free filter paper (retention: 2 µm) was performed.
 However, it should be mentioned that only negligible amounts of telluric acid remain in the filter paper.
 In the final step, the benzene is evaporated at room temperature with constant stirring, leaving behind a colorless, transparent, and highly viscous liquid.
-It should be noted here that benzene was deliberately chosen to enable evaporation at room temperature and normal pressure of the organic carrier medium.
+It should be noted that benzene was deliberately chosen to enable evaporation of the organic carrier medium at room temperature and atmospheric pressure.
 This ensures the removal of the by-product of the reaction (water) by means of an azeotropic distillation in a simple manner.
-It is clear that benzene could be replaced by toluene or cyclohexane if necessary, and that azeotropic distillation could be carried out using a rotary evaporator and low pressure at similarly low temperatures.
+It is clear that benzene could be replaced by toluene if necessary, and that azeotropic distillation could be carried out using a rotary evaporator and low pressure at similarly low temperatures.
 Despite its toxicity, benzene has proven to be ideal and particularly easy to use for the small quantities of Te-diols required here, while toluene should be preferred for later, larger applications.
@@
 The studies for this influence on the global light yield are presented in subsection 3.3.
+[H] Refer to caption
 Figure 1: Effective emission spectra for the unloaded Borexino scintillator and a sample loaded with 1% Te.
@@
 Within the scope of measurement uncertainties, no relevant deterioration in the absorbance of the samples could be observed.
-Loading with 0.125 % - 2 % of the Te-complex appears to have no relevant influence on the transmission of the Borexino mixture.
+Loading with 0.125% - 2% of the Te-complex shows no significant influence on the transmission of the Borexino mixture.
 For reasons of clarity, only the pure scintillator and the sample loaded with 1% are shown in Figure 2.
@@
 At the critical concentration (typ.
-<1<1 g/l) for the solvent-fluor combination, about half the maximum light yield is reached [12, 13].
+<1 g/l) for the solvent-fluor combination, about half the maximum light yield is reached [12, 13].
 For scintillation cocktails with effective energy transfers between solvent and fluor the critical concentration is lower, leading to higher light yields with lower amounts of fluor.
@@
 A 10.9 ml sample of LS is filled under protective N2 atmosphere into a highly reflective PTFE cell with 2 mm thin UV-transparent glass windows such that the entire volume is free of gas bubbles.
-Before and after the filling the sample as saturated with protective inert gas by extensively bubbling it with N2.
+Before and after the filling the sample was saturated with inert gas by extensively bubbling N2 through it.
 To protect the sample, the cell is sealed under moderate over-pressure.
-For the readout of the scintillation light, two conventional 1.13-inch PMTs (ETEL9128B) are coupled to the cell windows.
+For the readout of the scintillation light, two conventional 1.13 inch PMTs (ETEL9128B) are coupled to the cell windows.
 During the measurements the sample is irradiated by mono-energetic 662 keV gammas from a 137Cs source (activity: $\sim$370 kBq).
 The position of the source is thereby fixed with respect to the LS cell.
-The setup is triggered on a coincidence with a 1.5 inch $\times$ 1.5 inch LaBr3(Ce) detector provided by OST Photonics with 3.05% resolution at 662 keV.
+The setup is triggered by a coincidence with a $\times\,1.5 inch$ LaBr3(Ce) detector provided by OST Photonics with 3.05% resolution at 662 keV.
 In the offline analysis only events with a coincident signal in the 185 keV backscattering pseudo-peak of the crystal detector and thus a 477 keV energy deposition in the liquid scintillator are selected [15, 16].
+Figure 3: Schematic drawing of the LY setup.
+A small $1\times 1$ inch highly reflective PTFE liquid scintillator cell with thin UV-transparent glass windows is filled with the LS sample.
+Two PMTs are coupled to the cell.
+The LS is irradiated by mono-energetic gamma-quanta from a 137Cs source.
+To fix the scattering angle and by that the energy deposition of the gammas in the LS, the setup is triggered in coincidence with a LaBr3(Ce) detector.
+For further details on the setup and the measurement technique see also [15].
 The measured values for different loadings are shown in Table 2 and graphically depicted in Figure 4.
@@
 Each sample was measured 10 times on different days each time freshly filled into the previously cleaned empty cell.
-The variance of these measurements is interpreted as the influence of systematic effects like HV stability, minor temperature changes in the laboratory, efficiency of the oxygen removal from the LS and other sample preparation and cleaning related differences.
-By this study the relative systematic uncertainty on the LY of a given sample can be estimated as well as statistical errors.
-Figure 3: Schematic drawing of the LY setup.
-A small $1\times 1$ inch highly reflective PTFE liquid scintillator cell with thin UV-transparent glass windows is filled with the LS sample.
-Two PMTs are coupled to the cell.
-The LS is irradiated by monoenergetic gamma-quanta from a 137Cs source.
-To fix the scattering angle and by that the energy deposition of the gammas in the LS, the setup is triggered in coincidence with a LaBr3(Ce) detector.
-For further details on the setup and the measurement technique see also [15].
-Te Loading 0 % 0.125 % 0.25 % 0.5 % 1.0 % 2.0 %
+The variance of these measurements is interpreted as the influence of systematic effects like HV stability, minor temperature changes in the laboratory, efficiency of the oxygen removal from the LS and other sample preparation and cleaning-related differences.
+From this study, both the relative systematic uncertainty and the statistical error on the LY of a given sample can be estimated.
+Te-Loading 0% 0.125% 0.25% 0.5% 1.0% 2.0%
 Rel. LY [%] 100 $91.5\pm 3.42$ $82.3\pm 3.63$ $71.1\pm 3.50$ $61.9\pm 3.95$ $41.8\pm 4.01$
-Abs. LY [Photons/MeVee] 13600 $2444\pm 466 11$$1193\pm 494$ $9670\pm 476$ $8418\pm 538$ $5685\pm 546$
+[…]
 Table 2: The table shows the relative light yield values together with the Te concentration in the scintillator.
@@
 It should be mentioned, that light yields above 104 Photons for electron-like energy depositions are still realized with loadings well exceeding 0.25%.
-For loadings exceeding 1.0 % the chemical quenching by then Te-complex molecules severely reduces the light emission in this simple LS baesed on a single solvent (PC) and a single fluor (PPO).
+For loadings exceeding 1.0% the chemical quenching by then Te-complex molecules severely reduces the light emission in this simple LS based on a single solvent (PC) and a single fluor (PPO).
 3.4 Light Emission Time Profile
@@
 Right: Time profile experimental.
-In is possible to see: the cuvette on the bottom left, with inserted an optical fiber for the PMT characterization, the HL-PMt on the top left very close to the cell and on the right a neutral optical filter, behind which the $L$-PMT is mounted.
+It is possible to see: the cuvette on the bottom left, with inserted an optical fiber for the PMT characterization, the HL-PMT on the top left very close to the cell and on the right a neutral optical filter, behind which the $L$-PMT is mounted.
 The time response of the setup (IRF: Instrumental Response Function) is characterized in situ by the usage of a diode laser (EPL-405) with a typical pulse width of $\sim$70 ps.
@@
 3.4.2 Time Profiles for $\alpha$-interactions with the Liquid Scintillator
-Figure 6: Fluorescence time profile of the different mixtures normalised at maximum.
-On the left, the whole profile is shown, while on the right, a zoom in the interested region has been plotted.
-Due to the quenching effect on the light yield, the statistics for high concentrated Tellurium mixture is lower but enough to describe the fluorescence profile in the first hundreds of nanoseconds.
 For the time profile study presented in this publication a 244Cm alpha source was deployed into the LS vessel.
 After extensive flushing with nitrogen gas the data taking took place subsequently.
-The time profiles for different tellurium loadings of the LS are depicted in Figure 6 in logarithmic scale for both, the full DAQ window of 1600 ns (left plot) and a zoom in (right plot) on the first $\sim$ 100 ns of the light emission decay process.
+The time profiles for different tellurium-loadings of the LS are depicted in Figure 6 in logarithmic scale for both, the full DAQ window of 1600 ns (left plot) and a zoom in (right plot) on the first $\sim$100 ns of the light emission decay process.
+Figure 6: Fluorescence time profile of the different mixtures normalised at maximum.
+On the left, the whole profile is shown, while on the right, a zoom on the interested region has been plotted.
+Due to the quenching effect on the light yield, the statistics for high concentrated Tellurium mixture is lower but enough to describe the fluorescence profile in the first hundreds of nanoseconds.
 To better quantify changes in the time profile, a fit model based on four exponential decay functions convoluted with the IRF is applied to every dataset.
@@
 A more detailed description of the setup and the model are given in [21].
-$F_{fit}(t)=IRF(t-t_{0})\otimes N\sum_{i=1}^{4}\frac{A_{i}}{\tau_{i}-\tau_{r}}\left(e^{-\frac{t-t_{0}}{\tau_{i}}}-e^{-\frac{t-t_{0}}{\tau_{r}}}\right)$ (3.1)
-The results of all the fits are shown in table 3, in which it is possible to see the reduction of the decay time components for the all the components, while the fractions A are almost constant.
-Te Loading 0 % 0.125 % 0.25 % 0.5 % 1.0 % 2.0 %
-$A_{1}$ (%) $68.15\pm 0.15$ $9.17\pm 0.16 66$$6.75\pm 0.31$ $67.47\pm 0.19$ $8.10\pm 0.43 66$$6.86\pm 0.43$
-$A_{2}$ (%) $19\pm 0.18 14.14$$14\pm 0.14$ $14.68\pm 0.27$ $14.44\pm 0.21$ $14.68\pm 0.30$ $17.30\pm 0.30$
-$A_{3}$ (%) $9.60\pm 0.16$ $9.76\pm 0.11$ $10.12\pm 0.23$ $9.76\pm 0.17$ $9.27\pm 0.26$ $8.79\pm 0.26$
-$A_{4}$ (%) $8.20\pm 0.29$ $7.20\pm 0.24$ $8.80\pm 0.48$ $8.60\pm 0.33$ $8.00\pm 0.59$ $7.20\pm 0.59$
-$\tau_{1}$ (ns) $3.212\pm 0.001$ $2.789\pm 0.004$ $2.549\pm 0.008$ $2.341\pm 0.001$ $2.010\pm 0.001$ $1.501\pm 0.001$
-$\tau_{2}$ (ns) $15.05\pm 0.29$ $2.59\pm 0.20 11$$38\pm 0.35 11.$$1.48\pm 0.25$ $9.62\pm 0.25$ $6.59\pm 0.24$
-$\tau_{3}$ (ns) $74.10\pm 2.70$ $67.57\pm 1.12$ $65.36\pm 3.07$ $61.30\pm 2.40$ $55.90\pm 3.20$ $46.92\pm 3.20$
-$\tau_{4}$ (ns) $478.9\pm 30.3$ $470.0\pm 30.0$ $82.7\pm 40.2 44$$48.9\pm 30.4$ $385.0\pm 51.6$ $399.8\pm 51.6$
+$F_{\text{fit}}(t)=IRF(t-t_{0})\otimes N\sum_{i=1}^{4}\frac{A_{i}}{\tau_{i}-\tau_{r}}\left(e^{-\frac{t-t_{0}}{\tau_{i}}}-e^{-\frac{t-t_{0}}{\tau_{r}}}\right)$ (3.1)
+The results of all the fits are shown in Table 3, in which it is possible to see the reduction of the decay time components for all the components, while the fractions A are almost constant.
+Te-Loading 0% 0.125% 0.25% 0.5% 1.0% 2.0%
+$A_{1}$ [%] $68.15\pm 0.15$ $9.17\pm 0.16 66$$6.75\pm 0.31$ $67.47\pm 0.19$ $8.10\pm 0.43 66$$6.86\pm 0.43$
+$A_{2}$ [%] $19\pm 0.18 14.14$$14\pm 0.14$ $14.68\pm 0.27$ $14.44\pm 0.21$ $14.68\pm 0.30$ $17.30\pm 0.30$
+$A_{3}$ [%] $9.60\pm 0.16$ $9.76\pm 0.11$ $10.12\pm 0.23$ $9.76\pm 0.17$ $9.27\pm 0.26$ $8.79\pm 0.26$
+$A_{4}$ [%] $8.20\pm 0.29$ $7.20\pm 0.24$ $8.80\pm 0.48$ $8.60\pm 0.33$ $8.00\pm 0.59$ $7.20\pm 0.59$
+$\tau_{1}$ [ns] $3.212\pm 0.001$ $2.789\pm 0.004$ $2.549\pm 0.008$ $2.341\pm 0.001$ $2.010\pm 0.001$ $1.501\pm 0.001$
+$\tau_{2}$ [ns] $15.05\pm 0.29$ $2.59\pm 0.20 11$$38\pm 0.35 11.$$1.48\pm 0.25$ $9.62\pm 0.25$ $6.59\pm 0.24$
+$\tau_{3}$ [ns] $74.10\pm 2.70$ $67.57\pm 1.12$ $65.36\pm 3.07$ $61.30\pm 2.40$ $55.90\pm 3.20$ $46.92\pm 3.20$
+$\tau_{4}$ [ns] $478.9\pm 30.3$ $470.0\pm 30.0$ $82.7\pm 40.2 44$$48.9\pm 30.4$ $385.0\pm 51.6$ $399.8\pm 51.6$
 Table 3: Parameters of the $\alpha$ time profile models for different Te-loadings.
@@
 4 Conclusions
-Metal-loaded scintillators will play an important role in the search for neutrino-less double beta decay in the future.
+Metal-loaded liquid scintillators are expected to play an important role in the search for neutrino-less double beta decay in the future.
 Experiments such as SNO+ and KamLAND-Zen have already achieved initial success in this area.
@@
 The emission spectrum and transmission were virtually unaffected by even high doping factors.
-The decline in light yield is also moderate, meaning that even with 1% tellurium, approximately 8400 photons for 1 MeV energy deposition by an electron are emitted.
+The decline in light yield is also moderate, meaning that even with 1% tellurium, approximately 8,400 photons for 1 MeV energy deposition by an electron are emitted.
 The time profile of light emission by alpha particles was also studied.
-It was shown that the speed of scintillation increases significantly when the tellurium loading is increased.
+It was shown that the speed of scintillation increases significantly when the tellurium-loading is increased.
 This effect was expected, as the addition of large atoms or molecules to scintillators usually leads to chemical quenching effects that accelerate light emission while reducing the absolute light yield.
@@
 Acknowledgments
-This work benefited substantially from the support with seed funding from the Cluster of Excellence ORIGINS (DFG, German Research Foundation under Germany’s Excellence Strategy – EXC-2094 – 390783311).
-This work benefited from the support by the Detector Laboratory of the Cluster of Excellence PRISMA+.
-We are very grateful to the staff of the Detector Division around Dr.
-Quirin Weitzel.
-We would like to express our warmest thanks to Dr.
-Mihail Mondeshki (Chemistry Department, JGU Mainz) and his team for their support with chemical expertise, their unbiased interdisciplinary collaboration, and numerous discussions.
+We gratefully acknowledge seed funding from the Munich Cluster of Excellence ORIGINS (DFG, German Research Foundation under Germany’s Excellence Strategy – EXC-2094 – 390783311) and the support of the Detector Laboratory of the Mainz Cluster of Excellence PRISMA+.
+Special thanks go to the staff of the Detector Division, in particular Dr.
+Quirin Weitzel, for their practical assistance.
+We also warmly thank Dr.
+Mihail Mondeshki (Chemistry Department, JGU Mainz) and his team for their chemical expertise, constructive interdisciplinary collaboration, and numerous insightful discussions.
+We would like to thank the chemotechnical employee of the Mainz Institute for Physics, Joachim Strübig, and the glass apparatus manufacturer Rainer Jera.
 We are grateful to the University of Milan and to INFN of Milan which host the fluorescence time profile setup in the Liquid Scintillator laboratory.
 We would also thank Dr.
-Barbara Cacciania and Fatima Horuia for the help for this work.
+Barbara Caccianiga and Fatima Horuia for their help for this work.
 For countless detailed and inspiring discussions we would like to thank especially Prof.